The graph of x + 2y = -5 contains which point ?
A) (0,0)
B) (2,-1)
C) (-2,-1)
D) (-1,-2)

Answer by fcabanski(1391)

A point is (x,y). Plug in each choice to see if it works.


A. (0,0) 0 + 2*0 = 0 which is not -5: no.


B. (2,-1) 2+2*(-1) = 0 which is not -5: no.


C. (-2,1) -2+2(1) = 0 which is not -5: no


The answer must be D.


Checking -1 + 2(-2) = -5: yes!